The consumers problem is max u(C1, C2) s.t. C1 +5 < y1 C2 < y2 + (1+r)s C1 > 0, c220 Characterize the solution
Answer #1

max uccis (2) subject to at say C₂ = G + (1+4)S Cizo, zo atszy s=g-c. c2= Y q + (1+r) putting value ofs in ca a = Yg+ (1+r) (